OpenStudy (anonymous):

naturally occuring neon gas contains 90.92% neon-20, 0.257%neon-21, and 8.82% neon-22. the respective atomic masses are 19.992 amu, 20.994 amu, and 21.994 amu calculate the average atomic mass of neon

OpenStudy (lgg23):

\[A.M.=\frac{ (mass _{1}\times percent _{1})+(mass _{2}\times percent _{2}) }{ 100 }\] And since we can assume that we have 100% of a substance... \[A.M.=\frac{ (p_{Ne_{20}}\times m_{Ne_{20}})+(p_{Ne_{21}}\times m_{Ne_{21}})+(p_{Ne_{22}}\times m_{Ne_{22}}) }{ 100 }\] \[A.M.=\frac{ (90.92\times19.992)+(0.257\times20.994)+(8.82\times21.994) }{ 100 }\] The average atomic (AM) mass is 20.17.

OpenStudy (anonymous):

Thank you so much, it's the right answer and good explanation.